Bob is taking his son to look at colleges. The first college they plan to visit is 150 miles from their home. In the first hour they drive at a rate of 60 mph. If they want to reach their destination in 2.5 hours, what speed must they average for the remainder of their trip?
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

Bob is taking his son to look at colleges. The first college they plan to visit is 150 miles from their home. In the first hour they drive at a rate of 60 mph. If they want to reach their destination in 2.5 hours, what speed must they average for the remainder of their trip?
--------
rate = distance/time
-----
rate = (150-60*1)/(2.5-1) = 90/1.5 = 90/(3/2) = 60 mph
=================================
Cheers,
Stan H.
===================

Answer by mananth(16949)

let the distance traveled 1 hour at 60 mph
distance traveled = 60 miles
balance 90 miles and 1.5 hours
r=d/t
r=90/1.5
r=60 mph